Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day04


Crew 100B Log Book for March 1 2011

Photos of the day





Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Motivated and healthy

Experiment Status: All the medical, astrobiological, rover, and food experiments began today with unique results.

The meals today were very delicious, since it was a cooking day.

Productivity: We were very motivated today , since our health and safety officer, rover engineer and crew scientists continued to work on their experiments .

EVA’s: We completed four successful EVA’s such as one with drone, one for sample collection, one for MARS exploration and the last one for communication.

Plan for tomorrow: To start with special exercises with health and safety officer and myself undertaking the medical experiments again on all crew members. I shall also undertake my chemical-sensation experiment tomorrow.

Observations of the Sun via the radio telescope will also begin. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will occur. Soil samples will again be collected during three EVA’s tomorrow. One will be on foot, and three will use the ATV’s. Tomorrow will be a non-cooking day.


Commander's Report

Balwant Rai Reporting

Today we started our day with our continuing medical experiments as well as a new examination that checked every crew member’s medical health checkup.

The crew’s health will continue to be monitored by Jasdeep for the remainder of our stay here on Mars. Breakfast was same as we have fruits and cereals. Every member was checked for their blood-pressures and pulse rate. Afterwards, some of the crew started the first EVA of the day with creating an obstacle course for the rover, and continuing to get familiar with flying the drone. The rest of our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well.

Today our astronomer did calibrations for radio-telescope and she set up her own experiment to be performed. Tonight, she expects it to be a very clear night, so, she may observe new things. Three more EVA’s were performed, where soil samples were taken for the microbiology experiment and another one to look for fossils and meteorites. Later in the evening one more EVA was performed by few of our crew members to test communication skills. Today we had a problem with toilet. Our Hab engineers fixed it well with their skills.

Today is a cooking day, so our chief chef cooked something different, interesting and very fresh for lunch and dinner. Tomorrow, we will continue our medical and astrobiological experiments, as well as continue working on getting the rover experiments. Radio telescope working will also be taken into account.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.73

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.40

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): ?

Notes/Comments: We can't see de gasoline level in the Gasoline Tank but it isn't empty because we fill 2 jerry can yesterday.


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 19

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 13

· Potable Water Meter Reading:

Notes/Comments: we have some problems with toilets. We can't flush properly with bucketful. When we try to flush it, the water level goes down very slowly.


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: Tank 1 is full. Some drops go into Tank 2. Soon Tank 3 could be filled.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: no

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: no

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 73

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 62


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:

Clear


ANSART Matthieu

Science Report (Medical)

Jasdeep Kaur Reporting

Time: 0930-Blood pressure and pulse rate monitored by Zephyr technology and by manually.

Crew Members: Science Team: 6 crew members Balwant Rai, Rachel Dompnier , Quentin Bourges,Chrystal Latham,Matthieu Ansart and Jasdeep Kaur

Location: Hab

Objective: To monitor the blood pressure by different exercises and massages.

Observation: All crew members were monitored for their blood pressures. The blood pressures and heart rate was also monitored by Zephyr technology. Pulse was checked manually and electronically.

Results: As with respect to the principles to be followed on Mars, crew members cannot eat fresh fruits, juices, vegetables and other stuff. This needs very good maintainance of healthy body functions. This can be balanced by some exercise and massages. This is fourth day and all crew members exhibited excellent health status.

Lessons learned: Some of other yoga exercises are recommended and will be taught in coming days to be motivated.


Science Report (Medical/Cognition)

Balwant Rai Reporting

Cognitive performance: JBRCOGEuroMDRS

Crew cognitive performance is critical to the success of manned space missions. Crew performance continues to be dependent on numbers of factors such as cognitive health, physical health, and a healthy environment. Given the demanding scientific and engineering responsibilities of exploration crews, one of the leading indicators of overall crew performance is cognitive performance.

A sustained effort to track crews' cognitive performance along with potential factors of influence, both during actual space missions and as part of precursor activities on the ground, is therefore needed to expand our knowledge on how to optimize human performance in space.

To the best my knowledge no evaluation has been made on cognitive performance involved in Mars analogue EVA activity. Hence, this study is planned to find the effects of EVA on cognitive performance. Cognitive performance was measured by Cog State.

All members were selected for this study. The results showed significantly decrease in cognitive performance after EVA. So, Martians should participate in physical activities to have more good results.


Science Report (Astronomy)

Crystal Latham Reporting

Last night, I took the crew out for an evening observing session since it was quite clear outside. I gave a short astronomy lesson on the night sky. I pointed out some of the major winter constellations, like Orion, Cassiopeia, and Taurus. I also taught the crew how to find the North Star, Polaris, by using the last two stars in the scoop of the Big Dipper to point to it.

Since, it was a rather chilly night, the sky was extremely clear allowing us to see the Milky Way and many more stars than I was even used to seeing. We also managed to see a few shooting stars unexpectedly. I also gave an explanation on what a galaxy was as well as how astronomers find extrasolar planets and other solar systems, like our own. I also discussed star clusters like the Pleiades, and how stars form in nebulas like the Orion Nebula, which we were all able to see.

Today was spent working on a problem with the receiver for the radio telescope. After some consulting with more knowledgeable sources, I found that the problem was actually quite simple. It seems there are some loose solders, and this resulted in the inability of the receiver to stay on. Since I am not very good at soldering, I have remedied the problem by simply taping the volume knob of the receiver against the receiver box, which has allowed the receiver to stay on. Tonight, after the Sun sets, I will calibrate the telescope so that tomorrow I can begin observing the sun for solar storms and flares.

Weather permitting tonight, I plan to use the Musk Observatory to do some simple observing of some neat deep sky objects, like the Orion Nebula, so that some members of the crew can learn and understand more about astronomy.


EVA 8 Report

Rachel Dompnier Reporting

EVA #8, 9:15 to 10:15 am

Crew Members: Matthieu Ansart, Quentin Bourgès, Rachel Dompnier

Site Location: Behind the Hab

Transit Mode: Foot

Objectives: Testing the new recording application for the drone on iPad and improving with its piloting. Improving with the rover.

Results: We got some videos from the drone but after a few minutes it became uncontrollable. The rover worked better today.

Lessons Learned: We guess there is some kind of problem with the trim for the drone. We have to make some adjustment to offset the losses.


EVA 9 Report

Jasdeep Kaur reporting

EVA # 9, 12:15-13:00

Crew Members: Matthieu Ansart, Jasdeep Kaur, Balwant Rai

Site Location: N 38.39179, W 110.10832

Transit Mode: ATVs

Objectives: GeoMission to collect soil samples with designed scientific protocols to look for confirmation of microbes and extremophiles

Results: Sample was collected in a direction perpendicular to Hab in N-W direction. This area showed signs of characteristics of cracks, dried impressions as well as ground fissures with different patterns in an area that had a very peculiar blackish-gray color.

Lessons Learned: Helmets have to be safely and tightly secured before going on long ATV missions. Do not leave the elastic of helmet open.


EVA 10 Report

Matthieu Ansart Reporting

EVA # 10, 15:00-17:00

Crew Members: Matthieu Ansart, Balwant Rai, Quentin Bourges, Rachel Dompnier.

Site Location: N 38.41479, W 110.79053

Altitude : 1378m

Transit Mode: Foot

Objectives: Collect samples (meteorites or typical Martian rocks)

Results: Meteorite samples was collected behind a hill.

Lessons Learned: Space suits aren't convenient when it's sunny and warm. Two members of this expedition turn back to the MDRS because of warm conditions. When we go far from the Hab, it's better to use ATV .


EVA 11 Report

Quentin Bourges Reporting

EVA#11 5:25 pm to 6:05 pm

Crew members: Rachel Dompnier, Matthieu Ansart, Quentin Bourges

Site location: N 38.42590° W 110.78714°

Transit mode: ATVs

Objective: To get the GPS coordinates of the location where radio contact ceases with the Hab, in order to establish a clear area on the map where we are sure to be able to contact the Hab during EVAs.

Results: We traveled to the north of the Hab, following the path until we managed to reach the location where there was no contact with the Hab. The GPS of this location were recorded and the distance to this location was calculated to be 2.2 km.

Lessons learned: Now we have a clear idea of where we can go heading north until we lose communication with the Hab.


Chef's Report

Balwant Rai Reporting

1a) Today was

- A cooking day

- Chief Chef: Jasdeep Kaur

1) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

It was indeed the very special day and Jasdeep prepared a real food for crew members. Actually she dipped the dry-onions and tomatoes in morning, which were almost moist and could be used to cook on pan with olive oil, basil, oregano leaves to give a fresh taste. Also she added special flavors to the food. It was just fantastic. For dinner another Tofu-rice creation was prepared.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

1. Package of de-hydrated ham and pasta, vegetables, olive oil and bernaise sauce cooked as follows: in a bowl with a sheet of spaghetti below, sauce of tomatoes-onions cooked and prepared before added as a layer over it, pepperoni and ham layer added as another layer over and then cheddar cheese on top as a layer above served with bernaise sauce (lunch)

2. Packets of Tofu, rice, almond, groundnuts, peas, garlic, basil, bay leaves served with sauce(dinner)

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Nothing

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

No

5) Lessons learned (if any

Clean up the kitchen area before cooking 

6) Comments/questions for Kim and Jean, the food study investigators

No


Journalist Report

Crystal Latham Reporting

Its mission day four, here on Mars, and it was very fun filled. The day started with our now usual medical experiments, which are becoming a routine. Today another medical measurement was added to the mix that included taking the blood pressure and heart rate of each crew member, while at rest.

Three EVA’s took place today, and for those of you who don’t know, an EVA is an Extra-Vehicular Activity. The first EVA of the day was to continue testing the drone and rover’s maneuverability. Our second EVA collected more soil and rock samples, which have the potential to be real interesting finds. Our third and final EVA of the day had the objective of determining what the range of our radios work to. We found that at about 2 km, depending on the direction, signal is lost and mission control cannot be contacted. This has important implications for our exploration of the lunar surface.

Other highlights of our fourth day on Mars include getting the radio telescope receiver operational, as well as having a very large problem with the toilet. This problem has been fixed for the most part, but I know I speak for the crew when I say we would really love an Earth toilet out here on Mars.